RC Meeting, Sunday, September 5, 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meeting at 3:00 PM, (remote)

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Accelerator Status

  • As of yesterday on track to send beam into the hall today during swing for optics checks
  • As of yesterday Ops plans to have the hall swept mid-day --> checked with the PD this morning, also with Mike Aiken: sweeping may happen mid-day or later, whenever is convenient for Ops (I left it up to them)
  • Overnight: Ops had problems with beam steering at ARC 9; vertical dispersion too large
  • This morning: Mike Aiken and crew working on it: https://logbooks.jlab.org/entry/3901217
  • Ops HCO list: https://logbooks.jlab.org/entry/3901128

Hall Status

  • All checklists done by Friday, Walter O.K.-ed the hall transition to sweep and beyond
  • HMS Q2: https://logbooks.jlab.org/entry/3901108
  • Raster: https://logbooks.jlab.org/entry/3901099
  • Target camera: e-mailed Joe and Jack to tell them that there is an opportunity to fix it during the Monday access
  • Ops are still working on re-configuring their Hall C target screens: Sue confirmed she has everything she needs and said that Eric Forman will work on that; Ops proposed that, if target screens not ready before Monday day, one of the Hall C target experts can make a RED log entry stating that the Hall C target ladder is in Home, it will stay there and beam (type, current) is authorized to the Hall C dump; that's all Ops will need to send beam for tuning until Monday morning. After the Monday access all Hall C target screens (ours and their) must be fully updated and operational before receiving any beam
  • On out Target screen Nathan noticed that there is a discrepancy between the encoder value between the SHOULD and READBACK when target is in Home. Greg confirmed that the readback is as intended (-1. which is v.v.v. close to zero) but the should (-9000) doesn't make sense; we have been trying to resolve this... --> at 9:55 this morning I texted Chris K.
  • Monday morning access: Dave M. will condense H and D
  • Monday morning access: work on HMS Q2, Raster, Target screens if not resolved prior to that point

Notices in Effect for Shift Crew

Current Issues

  • HMS Q2 (won't prevent beam delivery until Monday)
  • Maybe raster (won't prevent beam delivery until Monday)
  • Hall C target camera (won't prevent beam delivery until Monday)
  • Ops needs to work on their screens to reflect the current Hall C target ladder configuration
  • We need to resolve the discrepancy between "should" and "readback" encoder value for target in Home